1. Market Equilibrium and Incidence of a Subsidy

In the market for hazelnuts (Q, measured in kilograms), monthly demand is given by Qd(p) = 280,000 – 20,000p, and monthly supply by QS(p) = 5,000p – 20,000, where p is the price of a kg of hazelnuts (in €). Suppose the government introduces a subsidy (s) of €5/kg.

a) Find the equilibrium price received by producers before (p*) and after (ps**) the subsidy.

[Hint: A subsidy is just a negative tax, so in equilibrium we must have: ps** = pd** + s, where pd** is the new equilibrium price paid by consumers after the introduction of the subsidy]
Answer #1

od 2 80000 - 20, 0p Pd = 280000 20,000 Pd = 140 20 ou 03 - 50000 - 200 + 20 DUO R 5000 5000 Before & body at eguilibrium PaaiPoice secrived by producer beter subsidy (p) - $12 After Subsidy = Pdts P = pd +5 Io + 4 = 14 - a 5000 20010 [ Subsidy (8) -